Mickley Corporation produces two products, Alpha6s and Zeta7s, which pass through two operations, Sintering and Finishing. Each of the products uses two raw materials—X442 and Y661. The company uses a standard cost system, with the following standards for each product (on a per unit basis):

Raw Material Standard Labor Time
Product X442 Y661 Sintering Finishing
Alpha6 2.0 kilos 1.5 liters 0.10 hours 0.80 hours
Zeta7 3.0 kilos 2.5 liters 0.45 hours 0.90 hours

Information relating to materials purchased and materials used in production during May follows:

Material Purchases Purchase Cost Standard
Price
Used in
Production
X442 14,800 kilos $66,600 $4.30 per kilo 9,300 kilos
Y661 15,800 liters $23,700 $1.60 per liter 13,800 liters

The following additional information is available:

  1. The company recognizes price variances when materials are purchased.
  2. The standard labor rate is $22.00 per hour in Sintering and $21.00 per hour in Finishing.

  3. During May, 1,220 direct labor-hours were worked in Sintering at a total labor cost of $28,060, and 2,930 direct labor-hours were worked in Finishing at a total labor cost of $67,390.

  4. Production during May was 2,300 Alpha6s and 1,600 Zeta7s.

Required:

1. Complete the standard cost card for each product, showing the standard cost of direct materials and direct labor.

Complete the standard cost card for each product, showing the standard cost of direct materials and direct labor. (Round your answers to 2 decimal places.)

Standard Quantity or Hours Standard Price or Rate Standard Cost
Alpha6:
Direct materials-X442 kilos per kilo
Direct materials-Y661 liters per liter
Direct labor-Sintering hours per hour
Direct labor-Finishing hours per hour
Total
Zeta7:
Direct materials-X442 kilos per kilo
Direct materials-Y661 liters per liter
Direct labor-Sintering hours per hour
Direct labor-Finishing hours per hour
Total

2. Compute the materials price and quantity variances for each material.

Compute the materials price and quantity variances for each material. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

Direct Materials Variances—Material X442:
Materials price variance
Materials quantity variance
Direct Materials Variances—Material Y661:
Materials price variance
Materials quantity variance

3. Compute the labor rate and efficiency variances for each operation.

Compute the labor rate and efficiency variances for each operation. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

Direct Labor Variances—Sintering:
Labor rate variance
Labor efficiency variance
Direct Labor Variances—Finishing:
Labor rate variance
Labor efficiency variance
Answer #1
Part-1:: Mickley Corporation
Standard Standard Standard
Quantity or Hours (A) Price or Rate (B) Cost (AXB)
Alpha6:
Direct materials-X442 2 kilos $4.30 per kilo $8.60
Direct materials-Y661 1.5 litres $1.60 per litre $2.40
Direct labor-Sintering 0.1 hours $22.00 per hour $2.20
Direct labor-Finishing 0.8 hours $21.00 per hour $16.80
tal $30.00
Zeta7:
Direct materials-X442 3 kilos $4.30 per kilo $12.90
Direct materials-Y661 2.5 litres $1.60 per litre $4.00
Direct labor-Sintering 0.45 hours $22.00 per hour $9.90
Direct labor-Finishing 0.9 hours $21.00 per hour $18.90
Totaal $45.70
Part-2
2. Material price variance = [SP-AP] x AQ purchased
Material X442 = [$4.30 - $4.50] x 14800 = $2960 U
Material Y661 = [$1.60 - $1.50] x 15800 = $1580 F
Material quantity variance = [SQ-AQ used] x SP
Material X442 = [{(2 x 2300) + (3 x 1600)} - 9300] x $4.30 = $430 F
Material Y661 = [{(1.5 x 2300) + (2.5 x 1600)} - 13800] x $1.60 = $10160 U
* X442 Actual Price per Kilos = 66600/14800=$4.50
**Y661, Actual Price Per Litre = 23700/15800=$1.50/ Litre
Direct material variance - Material X442
Material price variance 2960U
Material quantity variance 430F
Direct material variance - Material Y661
Material price variance 1580F
Material quantity variance 10160U
Part-3
3. Direct labor rate variance = [SR-AR] x AH worked
Sintering = [22 - 23] x 1220 = $2080 U
Finishing = [21 - 23] x 2930 = $5920 U
*Actual Rate/ Hour for Sintering :- $28060/1220=$23
**Actual Rate/ Hour for Finishing :- $67390/2930=$23
Direct labor efficiency variance = [SH-AH] x SR
Sintering = [{(0.10 x 2300) + (0.45 x 1600)} - 1220] x $22 = $5940 U
Finishing = [{0.80 x 2300) + (0.90 x 1600)} - 2930] x $21 = $7350 F
Direct labor variances - Sintering:
Labor rate variance 2080U
Labor efficiency variance 5940U
Direct labor variances - Finishing:
Labor rate variance 5920U
Labor efficiency variance 7350 F